# Polynomial Composition Activations: Unleashing the Dynamics of Large Language Models

### v1

@@ Abstract
Transformers have found extensive applications across various domains due to the powerful fitting capabilities. This success can be partially attributed to their inherent nonlinearity. Thus, in addition to the ReLU function employed in the original transformer architecture, researchers have explored alternative modules such as GeLU and SwishGLU to enhance nonlinearity and thereby augment representational capacity. In this paper, we propose a novel category of polynomial composition activations (PolyCom), designed to optimize the dynamics of transformers. Theoretically, we provide a comprehensive mathematical analysis of PolyCom, highlighting its enhanced expressivity and efficacy relative to other activation functions. Notably, we demonstrate that networks incorporating PolyCom achieve the optimal approximation rate, indicating that PolyCom networks require minimal parameters to approximate general smooth functions in Sobolev spaces. We conduct empirical experiments on the pre-training configurations of large language models (LLMs), including both dense and sparse architectures. By substituting conventional activation functions with PolyCom, we enable LLMs to capture higher-order interactions within the data, thus improving performance metrics in terms of accuracy and convergence rates. Extensive experimental results demonstrate the effectiveness of our method, showing substantial improvements over other activation functions. Code is available at https://github.com/BryceZhuo/PolyCom.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) have revolutionized the field of deep learning, facilitating unprecedented advancements in natural language processing (Radford et al., 2019), computer vision (Dosovitskiy et al., 2021), and beyond (Dong et al., 2018; Arnab et al., 2021). Characterized by their attention mechanisms, transformers excel at capturing intricate relationships within data, making them indispensable in contemporary machine learning applications. However, despite their widespread success, there remain opportunities for further refinement, particularly concerning the selection of activation functions. The activation function plays a crucial role in determining the output of each neuron within a neural network. Traditionally, simple nonlinearities such as Rectified Linear Unit (ReLU) (Nair & Hinton, 2010) and its variants (Hendrycks & Gimpel, 2016; So et al., 2021) have been favored due to their computational efficiency and ease of implementation. Although effective, these activation functions are inherently limited in their ability to model complex higher-order relationships within data. This limitation can be particularly restrictive in transformer architectures, where the ability to capture subtle and complex dependencies is essential.
@@
We posit that the integration of polynomial composition activations within transformer models can lead to enhanced performance in tasks requiring intricate data interpretation. To evaluate this hypothesis, we conducted comprehensive experiments on the pre-training configurations of large language models (LLMs), including both dense and sparse architectures. These evaluations were performed across various benchmarks, assessing the performance of transformers employing polynomial composition activations in comparison to those utilizing traditional activation functions. The results indicate that the proposed method not only improves model accuracy, but also accelerates convergence rates, thereby suggesting that polynomial composition activations provide a substantive advantage in deep learning applications.
@@
  • We propose a new activation function PolyCom which is a composition of the polynomial and other types of function. In particular, we introduce two instances of PolyCom: PolyReLU and PolyNorm, and details its integration into the transformer architecture.
@@ 2 Polynomial Composition Activation Function
PolyCom. The study of the polynomial activation function can be traced back to the seminal work of Hornik et al. (1989), which showed that neural networks with polynomial activation are not dense within the space of continuous functions. Additionally, empirical evidence has shown that deep neural networks employing pure polynomial activations tend to underperform (Trefethen, 2019). To overcome these limitations, we propose PolyCom, a novel composition of polynomial and other functions. Specifically, we explore two composition approaches:
@@
where $r\in{\mathbb{N}}$ denotes the order of PolyCom and $\rho$ represents an arbitrary functions such as ReLU, PReLU, Sigmoid, SiLU, or normalization. The key distinction between the two approaches lies in whether the function is composed before or after the power operation. It can be theoretically shown that both approaches have equivalent expressivity, provided that $\rho$ is a non-linear function. This is because polynomial terms are symmetric with respect to composition, allowing both Type I and Type II to approximate similar function classes. In other words, rearranging the order of $\rho$ and the polynomial powers does not affect the ability to approximate complex non-linear functions. In practice, we use a third-order PolyCom ($r=3$) with trainable coefficients $a_{i}$. For initialization, we set $a_{i}=1/r$ for $i=1,2,\dots,r$ and $a_{0}=0$.
For Type I PolyCom, we specifically consider a composition involving the ReLU function due to its simplicity, which we term PolyReLU. An $r$-order PolyReLU is defined as:
@@
[…]
For Type II PolyCom, we introduce PolyNorm, which normalizes the powers to ensure consistent magnitudes across terms:
@@
where ${\bm{x}}^{i}=[x_{1}^{i},x_{2}^{i},\cdots,x_{d}^{i}]^{\top}$ represents element-wise exponentiation, and $\|\cdot\|_{2}$ denotes the $L_{2}$ normalization.
Figure 2: Block diagrams of Transformer MLP blocks utilizing ReLU/GELU, SwiGLU, PolyReLU and PolyNorm. “FC” stands for Fully Connected layer. “$x^{i}$” represents the $i$-th power of the input tensor $x$, while “$a_{j}$” denotes the $j$-th element of the learnable weight vector $a$. “N” indicates a normalization operation.
@@
$\mathrm{FFN}_{\rho}({\bm{x}})=\rho({\bm{x}}W_{1})W_{2}$ (4)
where $\rho$ represents the activation function such as ReLU, GeLU, PolyReLU, or PolyNorm. We replace the traditional activation function with our proposed PolyCom variants to enhance model capacity and performance, as illustrated in Figure 2.
@@ 3 Theoretical Analysis
As discussed in Section 2, PolyReLU and PolyNorm have equivalent expressivity. To streamline the analysis, we focus solely on the theoretical properties of PolyReLU, specifically its expressivity and effectiveness. Additional, nonlinear activations such as GeLU and SwiGLU can be locally approximated by Taylor polynomials around the origin, which allows us to primarily compare PolyReLU with ReLU and polynomial activations. To avoid confusion, we refer to networks that use ReLU activations as ReLU networks, and those that use PolyReLU activations as PolyReLU networks.
@@ 3.1 Approximating ReLU Networks by PolyReLU
Proof of Lemma 1.
Thus, for any polynomial activation of order $r$,
@@ 3.2 Approximating PolyReLU with ReLU networks
For the activation $\mathrm{PolyReLU}(x)=\sum_{i=0}^{r}a_{i}\mathrm{ReLU}^{i}(x),x\in[-1,1]$ with $a_{i}\in[-1,1]$. Given any $\epsilon\in(0,1)$, there exists a ReLU network $f:[-1,1]\rightarrow[-1,1]$ with size $O(\ln^{2}(1/\epsilon))$, such that
@@
The proof is provided in the Appendix A. Lemma 2 establishes an upper bound on the size of a ReLU network needed to approximate a PolyReLU activation function. This result highlights that while ReLU networks can approximate PolyReLU activations, they require a significantly larger number of parameters.
@@
Let $g:[-1,1]^{d}\rightarrow[-1,1]$ be a PolyReLU network with depth $L$ and width $K$, and PolyReLU activation with order $r$ and Lipschitz constant $\alpha$. Suppose each neuron computes $x\mapsto\mathrm{PolyReLU}(a^{\top}x+b)$ with the pair $(a,b)$ satisfies $\|a\|_{1}+b\leq 1$ and $\mathrm{PolyReLU}:[-1,1]\rightarrow[-1,1]$ (a, b, and PolyReLU are possibly distinct across neuron). For any given $\epsilon\in(0,1)$, there exists a ReLU network $f:[-1,1]^{d}\rightarrow[-1,1]$ of size
@@
Conversely, the size of any ReLU network that approximates a PolyReLU network within tolerance $\epsilon$, must be at least
@@ 3.3 Approximation of General Smooth Function
${\mathcal{W}}^{n,\infty}\left([-1,1]^{d}\right)=\left\{f\in L^{\infty}\left([-1,1]^{d}\right)|\|f\|_{{\mathcal{W}}^{n,\infty}\left([-1,1]^{d}\right)}<\infty\right\},$ (12)
with the norm which defined as the following
@@
Theorem 3 indicates that PolyReLU networks can achieve an optimal approximation rate of $O(\epsilon^{-d/n})$. In contrast, previous works by Yarotsky (2017) demonstrated that ReLU networks require $O(\epsilon^{-d/n}\ln(1/\epsilon))$ parameters to achieve a similar approximation error. Similarly Boullé et al. (2020) showed that rational neural networks need $O(\epsilon^{-d/n}\ln(\ln(1/\epsilon)))$ parameters for the same task. Therefore, the approximation ability of PolyReLU networks is superior to that of both ReLU networks and rational networks. Furthermore, Theorem 4.2 in DeVore et al. (1989) shows that the total number of parameters required by neural networks to approximate functions in $F_{n,d}$ is $\Omega(\epsilon^{-d/n})$. Therefore, our PolyReLU networks achieve the optimal approximation rate in the context of Sobolev spaces.
@@ 4.1 Setup
Model Configuration. For the dense model, the transformer consists of 24 layers with hidden size $d_{model}=2048$ and 16 attention heads. In the MoE model, the transformer is composed of 16 layers, with a hidden size of $d_{model}=2048$, 16 attention heads, and 64 experts. To maintain a consistent number of trainable parameters across all activation functions, we adjust the intermediate size accordingly. Specifically, for SwiGLU, the intermediate size is set to two-thirds that of the other activations in all experiments.
@@
Hyperparameters. Unless otherwise specified, we use a third-order PolyCom by default and initialize the coefficients as $a_{i}=1/3$ for $i=1,2,3$ and set $a_{0}=0$. Model weights are randomly initialized. For optimization, we apply the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. All models are trained on sequences of 4096 tokens. For the dense model, we set the initial learning rate to 3e-4, decaying to 1.5e-5 using a cosine scheduler. The MoE model starts with a learning rate of 4e-4, also decaying according to a cosine schedule.
Evaluation To evaluate the performance of LLMs with PolyCom, we use a wide range of open benchmarks, including ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), CoQA (Reddy et al., 2019), Winogrande (Sakaguchi et al., 2021), MMLU (Hendrycks et al., 2021), BoolQ (Clark et al., 2019), COPA (Gordon et al., 2012), CSQA (Talmor et al., 2019), OBQA (Mihaylov et al., 2018), and SocialIQA (Sap et al., 2019). We utilize the LM Eval Harness (Gao et al., 2023) for standardized performance evaluation.
@@ 4.2 Results on Dense Model
Dowmstream Evaluation. Table 1 presents the training loss, validation perplexity, and downstream task accuracy (%) after processing 250 billion training tokens. The downstream tasks include ARC-Easy, ARC-Challenge, HellaSwag, PIQA, SciQ, and Winograde. More detailed results are provided in Appendix D. The results clearly demonstrate that the PolyCom family (PolyReLU and PolyNorm) outperforms the other activation functions. For instance, PolyNorm outperforms SwiGLU by an average margin of 1.21% across six downstream tasks. This underscores the expressivity and efficiency of PolyCom as an activation function in transformer models.
@@ 4.3 Results on MoE Model
Our experiments with MoE modes are based on OLMOE-1B-7B, which has 1 billion activate parameters and 7 billion total parameters (Muennighoff et al., 2024). Due to computational constraints, we compare only PolyNorm activation function, shown to perform best in dense models, with the widely used SwiGLU activation function, which is commonly employed in current LLM architectures.
Figure 3: Traning and validation loss on C4 and Wikipedia for MoE models with 200 billion training tokens. We compare models using SwiGLU and PolyNorm activation functions. PolyNorm demonstrates lower training and validation losses, indicating faster convergence.
Figure 4: Downstrean performance dynamics on HellaSwag, MMLU Var, ARC-Challenge, and SciQ for MoE models with 200 billion training tokens. Models with PolyNorm significantly outperform those with SwiGLU on downstream tasks.
@@
Dowmstream Evaluation. Table 2 presents the validation losses on 11 datasets. PolyNorm consistently achieves lower validation losses than SwiGLU across all datasets, with an average improvement of 0.02. In Table 3, we also observe that PolyNorm outperforms SwiGLU on 8 downstream tasks. These results highlight the superior performance of models using the PolyNorm activation function. Additional results can be found in Appendix E.
@@
Order of PolyCom. We first investigate the effect of different orders of PolyCom. We vary the order $r$ of PolyReLU in the set $\{2,3,4\}$ and plot the results in Figure 5(a). As seen, the convergence speed improves as the order increases. However, there is no noticeable difference between orders 3 and 4 in terms of convergence speed. Additionally, increasing the order can lead to computational overhead and overflow issues, particularly when using low-precision arithmetic. Based on these observations, we select $r=3$ as the default order for PolyCom in our experiments, balancing both performance and computational efficiency.
@@
Rank of Weights. To understand how PolyCom enhances model performance, we analyze the rank of the weights in each FNN layer of the transformer. We use the effective rank (Roy & Vetterli, 2007) to measure the effective dimensionality of weights and its definition is in Appendix C.2. Figure 6 shows that PolyReLU and PolyNorm result in higher weight ranks compared to other activation functions such as SwiGLU, GELU, and ReLU. A higher rank in the weight matrices indicates greater capacity for representing complex patterns in the data. These findings suggest that PolyCom improves the expressibility of transformers by allowing the FNN layers to better utilize their parameters, ultimately leading to better generalization on downstream tasks.

### v2

@@ Abstract
Transformers have found extensive applications across various domains due to their powerful fitting capabilities. This success can be partially attributed to their inherent nonlinearity. Thus, in addition to the ReLU function employed in the original transformer architecture, researchers have explored alternative modules such as GeLU and SwishGLU to enhance nonlinearity and thereby augment representational capacity. In this paper, we propose a novel category of polynomial composition activations (PolyCom), designed to optimize the dynamics of transformers. Theoretically, we provide a comprehensive mathematical analysis of PolyCom, highlighting its enhanced expressivity and efficacy relative to other activation functions. Notably, we demonstrate that networks incorporating PolyCom achieve the optimal approximation rate, indicating that PolyCom networks require minimal parameters to approximate general smooth functions in Sobolev spaces. We conduct empirical experiments on the pre-training configurations of large language models (LLMs), including both dense and sparse architectures. By substituting conventional activation functions with PolyCom, we enable LLMs to capture higher-order interactions within the data, thus improving performance metrics in terms of accuracy and convergence rates. Extensive experimental results demonstrate the effectiveness of our method, showing substantial improvements over other activation functions. Code is available at https://github.com/BryceZhuo/PolyCom.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) have revolutionized the field of deep learning, facilitating unprecedented advancements in natural language processing (Radford et al., 2019; Zeng et al., 2020; Li et al., 2024; Zuo et al., 2024), computer vision (Dosovitskiy et al., 2021; Wang et al., 2022), and beyond (Dong et al., 2018; Arnab et al., 2021; Wang et al., 2020). Characterized by their attention mechanisms, transformers excel at capturing intricate relationships within data, making them indispensable in contemporary machine learning applications. However, despite their widespread success, there remain opportunities for further refinement, particularly concerning the selection of activation functions. The activation function plays a crucial role in determining the output of each neuron within a neural network. Traditionally, simple nonlinearities such as Rectified Linear Unit (ReLU) (Nair & Hinton, 2010) and its variants (Hendrycks & Gimpel, 2016; Krotov & Hopfield, 2016; Li et al., 2019; So et al., 2021) have been favored due to their computational efficiency and ease of implementation. Although effective, these activation functions are inherently limited in their ability to model complex higher-order relationships within data. This limitation can be particularly restrictive in transformer architectures, where the ability to capture subtle and complex dependencies is essential.
@@
We posit that the integration of polynomial composition activations within transformer models can lead to enhanced performance in tasks requiring intricate data interpretation. To evaluate this hypothesis, we conducted comprehensive experiments on the pre-training configurations of large language models (LLMs), including both dense and sparse architectures. These evaluations were performed across various benchmarks, assessing the performance of transformers employing polynomial composition activations in comparison to those utilizing traditional activation functions. The results indicate that the proposed method not only improves the model accuracy but also accelerates convergence rates, thereby suggesting that polynomial composition activations provide a substantive advantage in deep learning applications.
@@
  • We propose a new activation function PolyCom which is a composition of the polynomial and other types of function. In particular, we introduce two instances of PolyCom: PolyReLU and PolyNorm, and detail its integration into the transformer architecture.
@@ 2 Polynomial Composition Activation Function
PolyCom. The study of the polynomial activation function can be traced back to the seminal work of Hornik et al. (1989), which showed that neural networks with polynomial activation are not dense within the space of continuous functions. Additionally, empirical evidence has shown that deep neural networks employing pure polynomial activations tend to underperform (Trefethen, 2019). To overcome these limitations, we propose PolyCom, a novel composition of polynomials and other functions. Specifically, we explore two composition approaches
@@
where $r\in{\mathbb{N}}$ denotes the order of PolyCom and $\rho$ represents an arbitrary function such as ReLU, PReLU, Sigmoid, SiLU, or normalization. The key distinction between the two approaches lies in whether the function is composed before or after the power operation. The distinction between the two approaches lies in whether the composition or the power is performed first. In the practical implementation of PolyCom, we use 3-order PolyCom ($r=3$) with trainable coefficients $a_{i}$. For initialization, we set $a_{i}=1/r$ for $i=1,2,\dots,r$ and $a_{0}=0$. Our experiments on large language models (LLMs) show that 3-order PolyCom can indeed achieve extraordinary performance.
For Type I PolyCom, we specifically consider a composition involving the ReLU function due to its simplicity, which we term PolyReLU. An $r$-order PolyReLU is defined as
@@
where $\mathrm{ReLU}^{i}(x)=\max\{x,0\}^{i}$. This formulation can be seen as an extension of both ReLU and square ReLU.
For Type II PolyCom, we introduce PolyNorm, which normalizes the powers to ensure consistent magnitudes across terms
@@
where ${\bm{x}}^{i}=[x_{1}^{i},x_{2}^{i},\cdots,x_{d}^{i}]^{\top}$ represents element-wise exponentiation, and $\|\cdot\|_{2}$ denotes the $L_{2}$ normalization. PolyNorm incorporates normalization operators to rescale different powers into a manageable range, thereby preventing excessively large or small values. This makes the training procedures more stable.
Figure 2: Block diagrams of Transformer MLP blocks utilizing ReLU/GELU, SwiGLU, PolyReLU and PolyNorm. “FC” stands for Fully Connected layer. “$x^{i}$” represents the $i$-th power of the input tensor $x$, “$a_{j}$” denotes the $j$-th element of the learnable weight vector $a$, “N” indicates a normalization operation.
@@
$\mathrm{FFN}_{\rho}({\bm{x}})=\rho({\bm{x}}W_{1})W_{2},$ (4)
where $\rho$ represents the activation function such as ReLU, GeLU, PolyReLU, and PolyNorm. We replace the traditional activation function with our proposed PolyCom variants to enhance model capacity and performance, as illustrated in Figure 2.
@@ 3 Theoretical Analysis
From Figure 1, one can see that the expressivity of PolyNorm is greater than or equal to that of PolyReLU. To streamline the analysis, we focus solely on the theoretical properties of PolyReLU, specifically its expressivity and effectiveness. Additional, nonlinear activations such as GeLU and SwiGLU can be locally approximated by Taylor polynomials around the origin, which allows us to primarily compare PolyReLU with ReLU and polynomial activations. To avoid confusion, we refer to networks that use ReLU activations as ReLU networks, and those that use PolyReLU activations as PolyReLU networks.
@@ 3.2 Approximating PolyReLU with ReLU networks
For the fixed positive integer $r$ and the activation $\mathrm{PolyReLU}(x)=\sum_{i=0}^{r}a_{i}\mathrm{ReLU}^{i}(x),x\in[-1,1]$ with $a_{i}\in[-1,1]$. Given any $\epsilon\in(0,1)$, there exists a ReLU network $f:[-1,1]\rightarrow[-1,1]$ with size $O(\ln^{2}(1/\epsilon))$, such that
@@
Lemma 2 establishes an upper bound on the size of a ReLU network needed to approximate a PolyReLU activation function. This result highlights that while ReLU networks can approximate PolyReLU activations, they require a significantly larger number of parameters.
@@
Let $g:[-1,1]^{d}\rightarrow[-1,1]$ be a PolyReLU network with depth $L$ and width $K$, and PolyReLU activation with order $r$ and Lipschitz constant $\alpha$. Suppose each neuron computes $x\mapsto\mathrm{PolyReLU}(a^{\top}x+b)$ with the pair $(a,b)$ satisfies $\|a\|_{1}+b\leq 1$ and $\mathrm{PolyReLU}:[-1,1]\rightarrow[-1,1]$ (a, b, and PolyReLU are possibly distinct across neurons). For any given $\epsilon\in(0,1)$, there exists a ReLU network $f:[-1,1]^{d}\rightarrow[-1,1]$ of size
@@
Conversely, there exists PolyReLU networks cannot be approximated within tolerance $\epsilon$ by any ReLU network with a size less than
@@ 3.3 Approximation of General Smooth Function
${\mathcal{W}}^{n,\infty}\left([-1,1]^{d}\right)=\left\{f\in L_{\infty}\left([-1,1]^{d}\right)|\|f\|_{{\mathcal{W}}^{n,\infty}\left([-1,1]^{d}\right)}<\infty\right\},$ (10)
with the norm which is defined as the following
@@
Theorem 3 indicates that PolyReLU networks can achieve an optimal approximation rate of $O(\epsilon^{-d/n})$. In contrast, previous works by Yarotsky (2017) demonstrated that ReLU networks require $O(\epsilon^{-d/n}\ln(1/\epsilon))$ parameters to achieve a similar approximation error. Similarly Boullé et al. (2020) showed that rational neural networks need $O(\epsilon^{-d/n}\ln(\ln(1/\epsilon)))$ parameters for the same task. Therefore, the approximation ability of PolyReLU networks is superior to that of both ReLU networks and rational networks. Furthermore, Theorem 4.2 in DeVore et al. (1989) shows that the total number of parameters required by neural networks to approximate functions in $F_{n,d}$ is $\Omega(\epsilon^{-d/n})$. Therefore, our PolyReLU networks achieve the optimal approximation rate in the context of Sobolev spaces. Additional disscution is included in Appendix B.
@@ 4.1 Setup
Model Configuration. For the dense model, the transformer consists of 24 layers with hidden size $d_{model}=2048$ and 16 attention heads. In the MoE model, the transformer is composed of 16 layers, with a hidden size of $d_{model}=2048$, 16 attention heads, and 64 experts. To maintain a consistent number of trainable parameters across all activation functions, we adjust the intermediate size accordingly. Specifically, for SwiGLU, the intermediate size is set to two-thirds that of the other activations in all experiments. More details can be found in Appendix E.
@@
Hyperparameters. Unless otherwise specified, we use a 3-order PolyCom by default and initialize the coefficients as $a_{i}=1/3$ for $i=1,2,3$ and set $a_{0}=0$. Model weights are randomly initialized. For optimization, we apply the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. All models are trained on sequences of 4096 tokens. For the dense model, we set the initial learning rate to 3e-4, decaying to 1.5e-5 using a cosine scheduler. The MoE model starts with a learning rate of 4e-4, also decaying according to a cosine schedule. We summarize the hyperparameters in Table 7.
Evaluation. To evaluate the performance of LLMs with PolyCom, we use a wide range of open benchmarks, including ARC-Easy (Clark et al., 2018), ARC-Challenge (ARC-C) (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), CoQA (Reddy et al., 2019), Winogrande (Sakaguchi et al., 2021), MMLU (Hendrycks et al., 2021), BoolQ (Clark et al., 2019), COPA (Gordon et al., 2012), CSQA (Talmor et al., 2019), OBQA (Mihaylov et al., 2018), and SocialIQA (Sap et al., 2019). We utilize the LM Eval Harness (Gao et al., 2023) for standardized performance evaluation.
@@ 4.2 Results on Dense Model
Downstream Evaluation. Table 1 presents the training loss, validation perplexity, and downstream task accuracy (%) after processing 250 billion training tokens. The downstream tasks include ARC-Easy, ARC-Challenge, HellaSwag, PIQA, SciQ, and Winograde. More detailed results are provided in Appendix I. The results clearly demonstrate that the PolyCom family (PolyReLU and PolyNorm) outperforms the other activation functions. For instance, PolyNorm outperforms SwiGLU by an average margin of 1.21% across six downstream tasks. This underscores the expressivity and efficiency of PolyCom as an activation function in transformer models.
@@ 4.3 Results on MoE Model
Our experiments with MoE modes are based on OLMOE-1B-7B, which has 1 billion activate parameters and 7 billion total parameters (Muennighoff et al., 2024). Due to computational constraints, we compare only the PolyNorm activation function, shown to perform best in dense models, with the widely used SwiGLU activation function, which is commonly employed in current LLM architectures.
Figure 3: Training and validation loss on C4 and Wikipedia for MoE models with 200 billion training tokens. We compare models using SwiGLU and PolyNorm activation functions. PolyNorm demonstrates lower training and validation losses, indicating faster convergence.
Figure 4: Dynamics of downstream performance on HellaSwag, MMLU Var, ARC-Challenge, and SciQ for MoE models with 200 billion training tokens. Models with PolyNorm significantly outperform those with SwiGLU on downstream tasks.
@@
Dowmstream Evaluation. Table 2 presents the validation losses on 11 datasets. PolyNorm consistently achieves lower validation losses than SwiGLU across all datasets, with an average improvement of 0.02. In Table 3, we also observe that PolyNorm outperforms SwiGLU on 8 downstream tasks. These results highlight the superior performance of models using the PolyNorm activation function. Additional results can be found in Appendix J.
@@
Order of PolyCom. We first investigate the effect of different orders of PolyCom. We vary the order $r$ of PolyReLU in the range $\{2,3,4\}$ and plot the results in Figure 5(a). As seen, the convergence speed improves as the order increases. However, there is no noticeable difference between orders 3 and 4 in terms of convergence speed. Additionally, increasing the order can lead to computational overhead and overflow issues, particularly when using low-precision arithmetic. Based on these observations, we select $r=3$ as the default order for PolyCom in our experiments, balancing both performance and computational efficiency.
@@
Rank of Weights. To understand how PolyCom enhances model performance, we analyze the rank of the weights in each FNN layer of the transformer. We use the effective rank (Roy & Vetterli, 2007) to measure the effective dimensionality of weights and its definition is in Appendix E.3. Figure 6 shows that PolyReLU and PolyNorm result in higher weight ranks compared to other activation functions such as SwiGLU, GELU, and ReLU. A higher rank in the weight matrices usually indicates a greater capacity for representing complex patterns in the data. These findings suggest that PolyCom improves the expressibility of transformers by allowing the FNN layers to better utilize their parameters, ultimately leading to better generalization on downstream tasks.
@@
Training Stability. Through extensive experiments, we find both PolyReLU and PolyNorm maintain a stable training process within transformer architectures. Our analysis indicates the combination of normalization operators in transformers and standard gradient clipping strategies effectively stabilizes training dynamics. We specifically design PolyNorm to address potential instability associated with BF16/FP16 precision formats, incorporating normalization operators that rescale powers to a manageable range, thus preventing excessively large or small values. This is particularly beneficial for FP16 training, as demonstrated in Appendix H. In contrast, PolyReLU lacks this normalization feature, which may result in stability issues in non-transformer architectures like ResNet. As shown in Figure 5(a), a 3rd order (our default setting) is sufficient. Based on these findings, we recommend the following configurations: (1) For transformer-based models, use either PolyNorm or PolyReLU. (2) For non-transformer models or those with lower stability, prefer PolyNorm.
Computational Overhead and Memory Footprint. We provide detailed analyses of the runtime and memory overhead for the proposed activation functions, including FLOPs ratios and memory consumption, which are included in Appendix F. Overall, after applying the gradient checkpointing technique, the overhead and memory footprint are acceptable, and there is negligible difference in the training budget required compared to the widely used SwiGLU.